# Fully Passive Quantum Conference Key Agreement

### v1

@@ Acknowledgements.
Appendix A Active CKA Security Proof [18]
@@ Appendix B Security
Using a four-user passive CKA as an example, one obtains estimates of yields, $Y_{m_{A}m_{B}m_{C}m_{D}}$, through performing decoy state analysis, with output photon numbers $m_{A},m_{B},m_{C},m_{D}$. While considering the local channel losses, one corrects the yield in the following way,
$Y_{n_{A}n_{B}n_{C}n_{D}}^{Z}=$ $\int_{-\Delta_{\phi}}^{\Delta_{\phi}}\int_{-\Delta_{\phi}}^{\Delta_{\phi}}...\int_{-\Delta_{\phi}}^{\Delta_{\phi}}\sum_{m_{A}=0}^{n_{A}}\sum_{m_{B}=0}^{n_{B}}\sum_{m_{C}=0}^{n_{C}}\sum_{m_{D}=0}^{n_{D}}P_{\phi_{A1},\phi_{A2}}\left(m_{A}\mid n_{A}\right)P_{\phi_{B1},\phi_{B2}}\left(m_{B}\mid n_{B}\right)$
$\quad P_{\phi_{C1},\phi_{C2}}\left(m_{C}\mid n_{C}\right)P_{\phi_{D1},\phi_{D2}}\left(m_{D}\mid n_{D}\right)\quad Y_{m_{A}m_{B}m_{C}m_{D}}^{Z}d\phi_{A1}d\phi_{A2}...d\phi_{D2}$ (9)
@@
Appendix C Branch Cutting Method
In passive CKA, in order to speed up the calculation, we introduce this method to not calculate some slice-choice combinations because those do not contribute meaningful key rate, or have zero key rate. These discarded combination satisfy at least one of the following criteria:
@@ Appendix C Branch Cutting Method
  • The final passive signal phases mismatch is large for each pair of parties. Specifically, for a pair of parties $i$ and $j$, we demand

### v2

@@ Acknowledgements.
Appendix A Decoy State Analysis - Two-Decoy
Decoy-state analysis is utilized to determine the upper bounds of the yields, denoted as $Y^{Z,j}_{m_{A}m_{B}m_{C}m_{D}}$, by applying linear programs to the set of equations, in the four-user case [34, 22]:
$G^{j}_{S_{ijkl}}=\sum_{m_{A},\ldots,m_{D}=0}^{\infty}Y_{m_{A},\ldots,m_{D}}^{j}\left\langle P\left(m_{A},m_{B},m_{C},m_{D}\right)\right\rangle_{S_{ijkl}}$ (8)
where $G^{j}_{S_{ijkl}}$ is a direct observable from the channel model, representing the conditional probability of a single-click event in a PE round, given the decoy state setting $S_{ijkl}$, $G^{j}_{S_{ijkl}}=\text{Pr}(\Omega_{j}|S_{ijkl})$ [22].
The average photon number distribution is given by
[…]
where $P_{\text{Poisson }}$ is the Poissonian distribution, and
[…]
$P_{\text{int }}\left(\mu_{A},\dots,\mu_{D}\right)=\frac{1}{\pi^{4}\sqrt{\mu_{A}\left(\mu_{\max}-\mu_{A}\right)}\sqrt{\mu_{B}\left(\mu_{\max}-\mu_{B}\right)}\sqrt{\mu_{C}\left(\mu_{\max}-\mu_{C}\right)}\sqrt{\mu_{D}\left(\mu_{\max}-\mu_{D}\right)}}$ (11)
is the intensity probability distribution [34]. The bounds obtained from the linear program can then be applied to Equation B1 for local loss correction and subsequently to calculate the phase error rate.
Appendix B Active CKA Key Rate[22]
@@ Appendix C Security
Taking a four-user fully passive CKA as an example, estimates of yields, denoted as $Y^{Z,j}_{m_{A}m_{B}m_{C}m_{D}}$, can be obtained by performing decoy state analysis. The output photon numbers are represented by $m_{A}$, $m_{B}$, $m_{C}$, and $m_{D}$ (see Appendix D), while $j$ indicates the single-photon event $\Omega_{j}$. Here, we explicitly use the superscript $Z$ to indicate that it is in the $Z$ basis for clarity. However, it will be omitted in subsequent equations. While considering the local channel losses, one corrects the yield in the following way,
$Y_{n_{A}n_{B}n_{C}n_{D}}^{j}=$ $\int_{-\Delta_{\phi}}^{\Delta_{\phi}}\int_{-\Delta_{\phi}}^{\Delta_{\phi}}...\int_{-\Delta_{\phi}}^{\Delta_{\phi}}\sum_{m_{A}=0}^{n_{A}}\sum_{m_{B}=0}^{n_{B}}\sum_{m_{C}=0}^{n_{C}}\sum_{m_{D}=0}^{n_{D}}P_{\phi_{A1},\phi_{A2}}\left(m_{A}\mid n_{A}\right)P_{\phi_{B1},\phi_{B2}}\left(m_{B}\mid n_{B}\right)$
$\quad P_{\phi_{C1},\phi_{C2}}\left(m_{C}\mid n_{C}\right)P_{\phi_{D1},\phi_{D2}}\left(m_{D}\mid n_{D}\right)\quad Y_{m_{A}m_{B}m_{C}m_{D}}^{j}d\phi_{A1}d\phi_{A2}...d\phi_{D2}$ (13)
@@
Appendix D Branch Cutting Method
In fully passive CKA, in order to speed up the calculation, we introduce this method to not calculate some slice-choice combinations because those do not contribute meaningful key rate, or have zero key rate. These discarded combination satisfy at least one of the following criteria:
@@ Appendix D Branch Cutting Method
  • The final fully passive signal phases mismatch is large for each pair of parties. Specifically, for a pair of parties $i$ and $j$, we demand